can you help me to solve this problem please:
SUzy has a piggy bank consisting of nickels and dimes. if there are 30 coins worth $1.90 how many nickels and dimes are in the bank?
thanks

Answer by stanbon(57387)

SUzy has a piggy bank consisting of nickels and dimes. if there are 30 coins worth $1.90 how many nickels and dimes are in the bank?
----------
Let number of nickels be "x" ;Value of these is 5x cents
---------
Number of dimes is "30-x" ; Value of these is 10(30-x)= 300-10x cents
------------
EQUATION:
value + value = 190 cents
5x + 300-10x = 190
-5x = -110
x = 22 (number of nickels)
30-x= 8 (number of dimes)
========
Cheers,
stan H.